# Raising the Bar of AI-generated Image Detection with CLIP

### v1

@@ Appendix A Additional ablation study
Appendix B Additional robustness analysis
@@
In Section 5 of the paper, in Tab. 6, we provide some results on the robustness of SoTA and proposed methods in the presence of image impairments. Due to lack of space, however, we show only some very synthetic results, averaged on generic post-processed images. Here, we analyze robustness in more detail as a function of the JPEG Quality Factor, which ranges from 100 to 60, and the resizing scale, which ranges from 125% to 25%. In addition, we consider also another compression method, WEBP, which is gaining popularity on social networks and has been rarely been used for augmentation. Images are from the SynthBuster dataset [4], generated by Stable Diffusion 2, SDXL, DALL·E 2, DALL·E 3, Midjourney, and Firefly. We show results (AUC) only for the most competitive (in terms of robustness) SoTA methods, Corvi and Ojha. For our method we consider again four variants similarly to the main paper: the versions using 1k or 10k real and fake reference images, and the corresponding versions, 1k+ and 10k+, where data have been augmented with randomly compressed/resized images, while keeping constant the total number of images.
@@
Ojha suffers some performance losses, up to 10 points, in the presence of compression, be it JPEG or WEBP, and a much more rapid decline with resizing, to the point of becoming useless in the presence of a 25% rescaling. Corvi was trained with strong augmentation and, in fact, is not affected at all by JPEG compression, while it presents acceptable losses with strong WEBP compression (not seen in training), and strong resizing. The most remarkable outcome of this experiment, however, is the impressive robustness of the CLIP-based detectors. Both the 1k+ and 10k+ versions, the ones with augmentation, are basically insensitive to compression, no matter JPEG or WEBP, and resizing. The versions without augmentation suffer some loss of performance but not nearly as dramatic as for the reference methods, and remain effective in all conditions.
@@
Appendix C Further generalization results
In this Section we extend our generalization analysis to additional data. In the main paper we included 6,000 real images and 18,$0$ fake images from 18 models, now we consider 8 more generators with 8,000 additional synthetic images for our test set (Tab. 7). The StyleGAN-T [63], eDiff-I [3], and GigaGAN [39] images are provided by the authors of the original papers. Instead, we ourselves generated images for Score-SDE [68], Stable Diff. [62], DiT [57], Deepfloyd-IF [44], Diff.ProjectedGAN [73], GALIP [70], and DDPM [34] using the pre-trained models available on-line. The other generated images come from three public datasets [71, 14, 4]. To avoid possible biases, performance metrics are evaluated using for each fake dataset images drawn from the real dataset used in training, indicated in the central columns of Table 1 of main paper and Table 7 of this supplemental.
AUC/Acc BigGAN EG3D
Diff. Proj.
GALIP
Taming
DALL·E
DDPM
Deepfloyd-IF
AVG
Method
GAN
Transf.
Mini
II stage
[71] Wang et al.
92.7 / 66.1
94.4 / 59.2
89.8 / 52.1
89.7 / 57.4
54.3 / 51.7
62.5 / 51.8
31.6 / 50.1
43.1 / 50.1
69.8 / 54.8
[9] PatchFor.
85.5 / 52.5
69.8 / 50.0
92.6 / 61.7
98.2 / 73.4
71.2 / 51.0
83.8 / 51.4
98.4 / 50.2
83.4 / 50.0
85.4 / 55.0
[33] Grag. et al.
98.7 / 94.2
98.9 / 93.8
100. / 72.5
96.2 / 79.5
90.3 / 76.4
83.4 / 62.5
49.7 / 45.1
71.2 / 61.3
86.0 / 73.2
[47] Liu et al.
94.7 / 81.3
99.0 / 86.3
99.5 / 84.9
94.3 / 78.4
95.4 / 78.9
98.4 / 88.1
22.8 / 49.6
97.4 / 86.8
87.7 / 79.3
[14] Corvi et al.
83.4 / 51.8
25.2 / 50.0
96.6 / 71.2
87.7 / 50.9
99.3 / 89.5
99.7 / 83.8
100. / 90.3
68.5 / 50.8
82.5 / 67.3
[69] LGrad
77.2 / 69.0
68.8 / 59.8
99.5 / 90.1
56.7 / 55.1
74.1 / 64.5
67.3 / 59.9
9.8 / 16.4
75.0 / 62.5
66.1 / 59.7
[56] Ojha et al.
99.6 / 96.4
92.6 / 82.5
97.4 / 75.2
98.6 / 89.9
94.1 / 84.8
97.1 / 84.9
77.7 / 68.2
60.8 / 50.0
89.7 / 79.0
[72] DIRE-1
99.8 / 95.3
50.1 / 50.0
49.8 / 51.6
100. / 96.7
73.1 / 72.4
99.7 / 96.5
23.1 / 50.0
99.4 / 95.8
74.4 / 76.0
[72] DIRE-2
98.6 / 82.4
46.1 / 50.0
50.2 / 51.4
99.3 / 81.8
77.0 / 66.2
98.7 / 81.8
14.0 / 49.8
95.6 / 80.5
72.4 / 68.0
Ours 1k
96.7 / 85.2
86.3 / 73.8
99.3 / 69.7
100. / 99.7
99.3 / 94.3
100. / 98.9
94.7 / 86.0
99.7 / 90.6
97.0 / 87.3
Ours 1k+
90.9 / 77.4
75.5 / 64.7
89.4 / 68.4
99.9 / 98.2
93.4 / 82.3
99.6 / 94.0
94.6 / 87.7
99.7 / 95.1
92.9 / 83.5
Ours 10k
98.2 / 85.9
90.8 / 81.4
99.8 / 90.1
100. / 99.9
99.6 / 94.3
100. / 99.1
94.9 / 88.3
99.8 / 88.1
97.9 / 90.9
Ours 10k+
91.7 / 75.0
80.7 / 72.5
88.9 / 74.6
100. / 97.8
92.5 / 77.2
99.8 / 93.5
95.9 / 79.5
99.8 / 92.8
93.7 / 82.8
@@ Appendix C Further generalization results
AUC/Acc BigGAN EG3D
Diff. Proj.
GALIP
Taming
DALL·E
DDPM
Deepfloyd-IF
AVG
Method
GAN
Transf.
Mini
II stage
[71] Wang et al.
82.5 / 55.5
84.7 / 52.2
80.8 / 53.2
92.2 / 59.2
66.2 / 50.6
66.7 / 50.5
69.6 / 50.2
47.9 / 49.9
73.8 / 52.7
[9] PatchFor.
58.7 / 50.3
60.3 / 50.0
55.6 / 50.0
72.3 / 50.3
50.4 / 50.1
71.0 / 50.1
82.9 / 51.0
69.7 / 50.0
65.1 / 50.2
[33] Grag. et al.
97.6 / 74.4
92.7 / 56.5
99.2 / 65.6
99.1 / 79.8
83.6 / 53.3
89.6 / 54.0
74.5 / 49.4
54.3 / 50.0
86.3 / 60.4
[47] Liu et al.
57.7 / 52.0
51.6 / 50.4
62.9 / 54.1
65.2 / 52.3
49.2 / 50.2
49.1 / 50.7
58.3 / 49.1
64.4 / 52.0
57.3 / 51.4
[14] Corvi et al.
77.3 / 53.0
64.5 / 59.7
93.3 / 59.8
87.4 / 51.5
94.4 / 76.7
97.4 / 74.8
74.7 / 59.6
70.4 / 50.9
82.4 / 60.7
[69] LGrad
53.5 / 51.6
53.2 / 51.1
56.9 / 51.1
57.0 / 50.3
51.3 / 51.1
41.7 / 49.1
56.0 / 49.7
38.7 / 48.8
51.0 / 50.4
[56] Ojha et al.
94.8 / 84.9
79.3 / 61.5
84.7 / 68.5
95.7 / 84.5
92.9 / 81.8
89.1 / 72.8
90.8 / 79.1
59.1 / 49.5
85.8 / 72.8
[72] DIRE-1
47.2 / 49.5
35.2 / 49.3
35.2 / 49.6
47.4 / 50.6
47.3 / 49.7
54.6 / 51.1
38.5 / 49.4
64.6 / 53.2
46.2 / 50.3
[72] DIRE-2
44.6 / 47.0
33.6 / 47.1
37.5 / 44.9
46.2 / 47.2
43.3 / 46.8
53.1 / 50.0
42.2 / 47.5
66.3 / 59.3
45.9 / 48.7
Ours 1k
84.0 / 74.8
64.4 / 59.5
75.3 / 68.3
98.5 / 93.0
85.3 / 75.0
96.5 / 89.0
76.7 / 66.6
90.6 / 80.1
83.9 / 75.8
Ours 1k+
88.9 / 78.2
60.0 / 54.6
70.5 / 60.9
99.8 / 98.3
86.5 / 77.3
98.9 / 93.6
96.1 / 84.7
99.4 / 96.2
87.5 / 80.5
Ours 10k
84.4 / 59.1
70.9 / 63.3
68.8 / 55.6
99.2 / 79.7
85.7 / 67.7
97.5 / 68.7
64.6 / 57.0
95.1 / 62.2
83.3 / 64.1
Ours 10k+
91.3 / 79.4
68.7 / 59.7
68.7 / 61.7
99.9 / 99.0
86.2 / 71.2
99.4 / 93.6
92.7 / 62.4
99.8 / 97.8
88.3 / 78.1
@@
In Tab. 8 we show the results in terms of AUC and Accuracy on such data for the four version of the proposed CLIP-based detector and for the SoTA methods described in the main paper (Section 5): Wang et al. [71], Grag et al. [33], PatchForensics [9], Liu et al. [47], LGrad [69], Corvi et al. [14], Ojha et al. [56], DIRE [72].
@@
Appendix D Attack scenario
In this section, we conduct some additional experiments on a specific attack scenario, where a malicious user takes real images and passes them through the autoencoder of a specific generative model. The objective is to insert the low-level traces typical of synthetic architectures in real images in order to fool forensic detectors and make everything appear as synthetic. To this end, we took 1000 real images from the RAISE dataset [16] and passed them through the autoencoders of three different versions of diffusion models: Stable Diffusion 1.4, Stable Diffusion 2.0 [62] and Stable Diffusion XL [58]. In Fig. 9 we show that such process insert exactly the same forensic clues in the real images and they appear indistinguishable from the generated ones, without altering their visual appearance (see Fig. 5 of the main paper).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: Top (from left to right): Fourier spectra of the noise residuals for synthetic images generated with Stable Diffusion 1.4, Stable Diffusion 2.0 and Stable Diffusion XL. Bottom: Fourier spectra of the noise residuals for the real images that passed through the corresponding autoencoder of the synthetic generators. We can observe that this operation has inserted in the real images exactly the same artifacts present in the synthetic ones.
In Table 10 we report results in terms of accuracy before and after the attack for our method and the best competitors. Before the attack, the detector must discriminate between real and generated images while, after the attack, it must discriminate between laundered real images and generated ones. We notice that before the attack the method proposed by Corvi et al. can achieve almost perfect results. However, it is very sensitive to the insertion of low-level traces and provides totally wrong results on the attacked data. For Ojha et al the performance also worsen, with an average 7% loss. Instead, for our method the Accuracy is almost unaffected by the attack, with an average loss below 1.3%.
Based on these results, it appears that the proposed CLIP-based detector does not rely on the same low-level traces exploited by most of the current detectors. This ensures higher robustness to common attacks and opens the way to the use of suitable fusion strategies. However, there is much room for further analysis, to begin with identifying which traces our detector is actually exploiting.
Accuracy before the attack after the attack
Method
St.Diff.
St.Diff.
SDXL
St.Diff.
St.Diff.
SDXL
1.4
2.0
1.4
2.0
Corvi et al.
100.
99.2
99.7
50.0
49.2
53.7
Ojha et al.
69.9
75.6
68.0
59.0
69.0
64.5
Ours 1k
63.6
64.3
60.2
63.6
63.8
59.2
Ours 1k+
76.8
77.0
75.2
77.0
75.5
72.6
Ours 10k
65.8
60.5
54.4
65.9
60.2
54.0
Ours 10k+
78.8
72.8
70.0
79.1
71.5
68.7
Table 10: Results in terms of Accuracy before and after the laundering attack on real images. Before the attack all methods achieve a very good performance. After the attack the performance of our method is almost unchanged, while significant losses are observed for the best reference methods, especially for Corvi et al. that goes from perfect detection to random guess.
Appendix E Few-shot analysis in the wild
In this section we present an experiment in a few-shot scenario. The idea is to explore the ability of our detector to work with very limited data and adapt to a situation where only a few real and synthetic examples are available. We downloaded real and synthetic images from Twitter and considered three different generators DALL·E 3, Midjourney and Firefly (some examples can be seen in Fig. 10). To understand which generator was used to create a specific image, we relied on tags and annotations present on the social network.
In our few-shot scenario, we take $10$ real images and $10$ generated images as examples from a specific model and test on all the others images (note that we present average results on 1000 runs). Results are reported in Tab. 11 in terms of AUC and Accuracy. The availability of just 10+10 images of the target data provides an impressive performance boost with respect to the same method trained on a larger dataset, but not aligned with the test data (real images come from COCO, synthetic images from Latent diffusion models). It is important to underline that this is a realistic scenario, in which one is called to decide on images generated by a new method and a few sample images are available as a support. We believe that this can represent an interesting direction for the application in the wild or to easily adapt a detector to more challenging situations where some prior information is available.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Some examples of images downloaded from the social network Twitter. From left to right: a real image, synthetic images from DALL·E 3 [6], Firefly [30], Midjourney [52].
AUC/Acc Twitter
Method
DALL·E3
Midjourney
FireFly
Average
Ours 1k
71.9 / 51.1
70.3 / 54.0
77.4 / 52.4
73.2 / 52.5
Ours 1k+
83.6 / 72.3
75.5 / 67.3
89.6 / 80.5
82.9 / 73.4
Ours 10k
68.2 / 50.5
72.0 / 51.5
70.7 / 51.5
70.3 / 51.2
Ours 10k+
67.0 / 60.0
78.3 / 71.0
74.3 / 68.5
73.2 / 66.5
Ours Few-Shot
97.7 / 90.6
91.9 / 82.5
96.6 / 88.4
95.4 / 87.2
Table 11: We show results in a few-shot experiment, where we suppose to know in advance 10+10 real/fake images from a specific generator and compare them with our basic approach trained on a larger set of real/fake images from misaligned data. As can be seen, even with just 10 images it is possible to adapt the model and obtain a large improvement.

### v2

@@ Appendix C AUC vs Accuracy
Appendix D Additional robustness analysis
@@
In Section 5 of the main paper we provide some results on the robustness of SoTA and proposed methods in the presence of image impairments. Due to lack of space, however, we show only some very synthetic results, averaged on all kinds of post-processed images. In Fig. 8, we analyze robustness in more detail as a function of JPEG Quality Factor, ranging from 100 to 60, and resizing scale, going from 125% to 25%. In addition, we consider also another compression method, WebP, which is gaining popularity on social networks but has been rarely considered in experimental analyses. Images are from the SynthBuster dataset [4], generated by Stable Diffusion 2, SDXL, DALL·E 2, DALL·E 3, Midjourney, and Firefly. We show results (AUC) only for the SoTA methods that proved most competitive in terms of robustness, Corvi and Ojha. For our method we consider again the versions with 1k or 10k real and fake reference images, and the corresponding versions, 1k+ and 10k+, with the same number of images including random compression and resizing.
@@
Ojha suffers some performance losses, up to 10 points, in the presence of compression, be it JPEG or WebP, and a much more rapid decline with resizing, to the point of becoming useless in the presence of a 25% rescaling. Corvi was trained with strong augmentation and, in fact, is not affected at all by JPEG compression, while it presents acceptable losses with strong WebP compression (not seen in training), and strong resizing. The most remarkable outcome of this experiment, however, is the impressive robustness of the CLIP-based detectors. Both the 1k+ and 10k+ versions, those with augmentation, are basically insensitive to compression, no matter JPEG or WebP, and resizing. The versions without augmentation suffer some loss of performance but not nearly as dramatic as for the reference methods, and remain effective in all conditions.
AUC/Acc Social network: X
Method DALL·E3 Midjourney FireFly Average
Wang et al. 22.5 / 50.0 36.7 / 50.0 65.3 / 50.0 41.5 / 50.0
PatchFor. 54.5 / 50.0 59.9 / 50.0 46.5 / 50.0 53.6 / 50.0
Grag. et al. 44.7 / 49.6 58.4 / 49.7 80.5 / 49.9 61.2 / 49.7
Mand. et al. 42.9 / 49.3 53.7 / 50.6 52.7 / 50.1 49.8 / 50.0
Liu et al. 34.5 / 49.7 38.7 / 49.7 54.9 / 49.9 42.7 / 49.8
Corvi et al. 89.3 / 86.8 98.3 / 90.8 79.1 / 52.1 88.9 / 76.6
LGrad 33.7 / 39.0 37.9 / 41.4 61.4 / 58.4 44.3 / 46.3
Ojha et al. 34.4 / 45.6 32.2 / 46.2 73.0 / 65.7 46.5 / 52.5
DIRE-1 54.3 / 52.0 60.7 / 54.3 55.1 / 51.1 56.7 / 52.5
DIRE-2 60.0 / 55.0 58.6 / 53.9 47.7 / 46.5 55.4 / 51.8
NPR 43.8 / 41.7 46.2 / 44.9 65.4 / 62.0 51.8 / 49.5
Ours 1k 69.4 / 50.9 70.2 / 53.3 69.3 / 51.4 69.6 / 51.9
Ours 1k+ 82.1 / 73.3 82.8 / 74.5 83.8 / 74.9 82.9 / 74.3
Ours 10k 65.8 / 50.0 69.3 / 51.7 70.7 / 50.9 68.6 / 50.9
Ours 10k+ 78.7 / 63.7 82.2 / 69.6 82.7 / 69.6 81.2 / 67.6
Ours fusion 90.0 / 85.9 97.5 / 93.8 77.8 / 57.0 88.4 / 78.9
Ours fusion+ 94.7 / 87.3 97.9 / 92.1 85.4 / 74.2 92.6 / 84.5
Ours Few-Shot 98.5 / 92.5 95.8 / 87.4 92.5 / 83.5 95.6 / 87.8
Table 10: Results in terms of AUC and accuracy of proposal and SOTA methods on real and synthetic images download from X. The last row reports a few-shot experiment, where we suppose to know in advance 10+10 real/fake images from a specific generator.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Some examples of images downloaded from the social network X. From left to right: a real image, synthetic images from DALL·E 3 [6], Firefly [31], Midjourney [53].
Appendix E Few-shot analysis in the wild
In this section we present an experiment in a few-shot scenario. The idea is to explore the ability of our detector to work with very limited data and adapt to a situation where only a few real and synthetic examples are available. We downloaded 500 real images and 1,500 synthetic images from X from three different generators DALL·E 3, Midjourney and Firefly (some examples can be seen in Fig. 9). To understand which generator was used to create a specific image, we relied on tags and annotations present on the social network.
In our few-shot scenario, we take $10$ real images and $10$ generated images as examples from a specific model and test on all the others images (note that we present average results on 1,000 runs). Results are reported in Tab. 10 in terms of AUC and accuracy. The availability of just 10+10 images of the target data provides an impressive performance boost with respect to the same method trained on a dataset much larger but not aligned with the test data (real images come from COCO, synthetic images from Latent diffusion models). It is important to underline that this is a realistic scenario, in which one is called to decide on images generated by a new method and a few sample images are available as a support. We believe that this can represent an interesting direction for the application in the wild or to easily adapt a detector to more challenging situations where some prior information is available.
To complete our analysis on this dataset we include in Tab. 10 the comparison with SoTA methods and our original proposal. We can observe that the performance of all the methods degrade which highlights the increased difficulty to handle a realistic scenario over the web. The best performance can be obtained by using our fusion approach that take the best of the low-level and high-level features and is able to achieve on average more than 90% in terms of AUC which is comparable with the few-shot analysis where some prior knowledge is available.
@@ Appendix E Few-shot analysis in the wild
AUC/Acc BigGAN EG3D Diff. Proj. GALIP Taming DALL·E DDPM Deepfloyd-IF AVG
Method GAN Transf. Mini II stage
Wang et al. 92.7 / 66.1 94.4 / 59.2 89.8 / 52.1 89.7 / 57.4 54.3 / 51.7 62.5 / 51.8 31.6 / 50.1 43.1 / 50.1 69.8 / 54.8
PatchFor. 85.5 / 52.5 69.8 / 50.0 92.6 / 61.7 98.2 / 73.4 71.2 / 51.0 83.8 / 51.4 98.4 / 50.2 83.4 / 50.0 85.4 / 55.0
Grag. et al. 98.7 / 94.2 98.9 / 93.8 100. / 72.5 96.2 / 79.5 90.3 / 76.4 83.4 / 62.5 49.7 / 45.1 71.2 / 61.3 86.0 / 73.2
Mand. et al. 92.2 / 83.0 100. / 99.8 64.5 / 49.9 77.6 / 59.1 91.8 / 84.1 83.6 / 69.8 99.9 / 97.4 49.2 / 48.9 82.3 / 74.0
Liu et al. 94.7 / 81.3 99.0 / 86.3 99.5 / 84.9 94.3 / 78.4 95.4 / 78.9 98.4 / 88.1 22.8 / 49.6 97.4 / 86.8 87.7 / 79.3
Corvi et al. 83.4 / 51.8 25.2 / 50.0 96.6 / 71.2 87.7 / 50.9 99.3 / 89.5 99.7 / 83.8 100. / 90.3 68.5 / 50.8 82.5 / 67.3
LGrad 77.2 / 69.0 68.8 / 59.8 99.5 / 90.1 56.7 / 55.1 74.1 / 64.5 67.3 / 59.9  9.8 / 16.4 75.0 / 62.5 66.1 / 59.7
Ojha et al. 99.6 / 96.4 92.6 / 82.5 97.4 / 75.2 98.6 / 89.9 94.1 / 84.8 97.1 / 84.9 77.7 / 68.2 60.8 / 50.0 89.7 / 79.0
DIRE-1 99.8 / 95.3 50.1 / 50.0 49.8 / 51.6 100. / 96.7 73.1 / 72.4 99.7 / 96.5 23.1 / 50.0 99.4 / 95.8 74.4 / 76.0
DIRE-2 98.6 / 82.4 46.1 / 50.0 50.2 / 51.4 99.3 / 81.8 77.0 / 66.2 98.7 / 81.8 14.0 / 49.8 95.6 / 80.5 72.4 / 68.0
NPR 86.8 / 77.2 53.3 / 57.5 100. / 99.2 90.7 / 77.6 80.2 / 69.2 79.0 / 73.3 92.4 / 61.6 90.9 / 76.8 84.2 / 74.1
Ours 1k 96.9 / 86.1 87.0 / 73.0 99.3 / 70.2 100. / 99.7 99.4 / 95.0 100. / 99.1 95.1 / 86.7 99.7 / 90.7 97.2 / 87.6
Ours 1k+ 92.0 / 80.0 76.4 / 65.5 89.6 / 67.7 99.9 / 98.6 93.5 / 83.0 99.6 / 95.0 94.9 / 87.4 99.7 / 96.0 93.2 / 84.1
Ours 10k 98.2 / 87.0 87.7 / 63.2 99.7 / 87.4 100. / 99.8 99.7 / 97.4 100. / 99.4 93.9 / 61.9 99.8 / 85.9 97.4 / 85.3
Ours 10k+ 93.1 / 77.5 79.9 / 58.8 92.5 / 76.0 100. / 97.0 94.9 / 79.5 99.9 / 94.4 98.1 / 91.5 99.8 / 92.2 94.8 / 83.3
@@
AUC/Acc BigGAN EG3D Diff. Proj. GALIP Taming DALL·E DDPM Deepfloyd-IF AVG
Method GAN Transf. Mini II stage
Wang et al. 82.5 / 55.5 84.7 / 52.2 80.8 / 53.2 92.2 / 59.2 66.2 / 50.6 66.7 / 50.5 69.6 / 50.2 47.9 / 49.9 73.8 / 52.7
PatchFor. 58.7 / 50.3 60.3 / 50.0 55.6 / 50.0 72.3 / 50.3 50.4 / 50.1 71.0 / 50.1 82.9 / 51.0 69.7 / 50.0 65.1 / 50.2
Grag. et al. 97.6 / 74.4 92.7 / 56.5 99.2 / 65.6 99.1 / 79.8 83.6 / 53.3 89.6 / 54.0 74.5 / 49.4 54.3 / 50.0 86.3 / 60.4
Mand. et al. 71.4 / 55.5 85.8 / 62.1 60.3 / 49.6 78.1 / 57.3 82.3 / 59.7 68.2 / 57.4 61.1 / 58.6 49.0 / 49.8 69.5 / 56.3
Liu et al. 57.7 / 52.0 51.6 / 50.4 62.9 / 54.1 65.2 / 52.3 49.2 / 50.2 49.1 / 50.7 58.3 / 49.1 64.4 / 52.0 57.3 / 51.4
Corvi et al. 77.3 / 53.0 64.5 / 59.7 93.3 / 59.8 87.4 / 51.5 94.4 / 76.7 97.4 / 74.8 74.7 / 59.6 70.4 / 50.9 82.4 / 60.7
LGrad 53.5 / 51.6 53.2 / 51.1 56.9 / 51.1 57.0 / 50.3 51.3 / 51.1 41.7 / 49.1 56.0 / 49.7 38.7 / 48.8 51.0 / 50.4
Ojha et al. 94.8 / 84.9 79.3 / 61.5 84.7 / 68.5 95.7 / 84.5 92.9 / 81.8 89.1 / 72.8 90.8 / 79.1 59.1 / 49.5 85.8 / 72.8
DIRE-1 47.2 / 49.5 35.2 / 49.3 35.2 / 49.6 47.4 / 50.6 47.3 / 49.7 54.6 / 51.1 38.5 / 49.4 64.6 / 53.2 46.2 / 50.3
DIRE-2 44.6 / 47.0 33.6 / 47.1 37.5 / 44.9 46.2 / 47.2 43.3 / 46.8 53.1 / 50.0 42.2 / 47.5 66.3 / 59.3 45.9 / 48.7
NPR 52.6 / 50.0 40.2 / 49.8 48.8 / 50.3 57.8 / 50.6 46.2 / 49.8 51.7 / 49.6 54.5 / 50.5 51.3 / 48.9 50.4 / 49.9
Ours 1k 84.6 / 75.8 64.8 / 60.3 73.4 / 66.7 98.3 / 92.6 85.1 / 75.8 96.2 / 88.9 76.1 / 67.7 90.0 / 79.9 83.6 / 76.0
Ours 1k+ 89.6 / 79.7 60.5 / 55.1 69.6 / 61.0 99.9 / 98.4 86.9 / 78.0 98.9 / 93.7 96.4 / 84.8 99.5 / 96.5 87.7 / 80.9
Ours 10k 84.1 / 67.0 64.3 / 54.9 73.7 / 62.9 98.1 / 83.9 86.2 / 74.0 96.8 / 77.8 70.9 / 60.7 91.3 / 67.5 83.1 / 68.6
Ours 10k+ 92.4 / 82.4 63.1 / 55.5 74.1 / 63.9 99.9 / 98.9 88.1 / 78.2 99.4 / 94.9 96.1 / 87.1 99.7 / 97.0 89.1 / 82.2
@@
Appendix F Further generalization results